# Neural reproducing kernel Banach spaces and representer theorems for deep networks

### v1

@@ 5. Conclusion
Studying function spaces defined by neural networks provides a natural way to understand their properties. Recently reproducing kernel Banach spaces have emerged has a useful concept to study shallow networks.
@@
In this paper, we take a step towards more complex architectures considering deep networks. We allow for a wide class of activation functions and remove unnecessary low rank constraints. Our main contribution is defining a class of neural RKBS obtained composing vector valued RKBS and deriving corresponding representer theorems borrowing ideas from [18].
@@
Future developments include considering more structured architectures, for example convolutional networks, as well as investigating the statistical and computational properties of neural RKBS
@@ Appendix A Sparse solutions to finite-dimensional variational problems
Let $Q$ be a convex subset of a locally convex space. A point $q\in Q$ is called extremal if $Q\setminus\{q\}$ is convex, that is, there do not exist $p,r\in Q$, $p\neq r$, such that $q=tp+(1-t)r$ for some $t\in(0,1)$. We denote the set of extremal points of $Q$ by $\operatorname{Ext}(Q)$.
@@
A.2 is a simplified version of [7, Theorem 3.3], where $G$ is only assumed to be a seminorm.
In view of A.2, in order to determine the form of sparse solutions we need to characterize the set of extremal point of the unit ball in $U$. In our construction of integral RKBS, $U$ is the space of vector measures with total variation norm. The following result provides the desired characterization for our case. It appears in [29] assuming that $\Theta$ is compact, but the proof works analogously when $\Theta$ is locally compact. We report the proof for the reader’s convenience.
Lemma A.3 ([29, Theorem 2]).
@@
Let us denote $E=\{y\cdot\delta_{\theta}:y\in\operatorname{Ext}(B_{\mathcal{Y}}(1)),\theta\in\Theta\}$. We start showing that $\operatorname{Ext}(B_{\mathcal{M}}(1))\subseteq E$. Suppose that $\mu\in\operatorname{Ext}(B_{\mathcal{M}}(1))$ but $\mu\neq y\cdot\delta_{\theta}$ for any $y\in\operatorname{Ext}(B_{\mathcal{Y}}(1))$ and $\theta\in\Theta$. Then $|\mu|\neq\delta_{\theta}$ for any $\theta\in\Theta$, and there is $A\in\mathcal{B}(\Theta)$ such that $10<|\mu|(A)<$ 1. Denote by $\chi_{A}$ the indicator function on $A$. Then, setting $t=|\mu|(A)$, $\mu_{1}=\mu\chi_{A}/t$ and $\mu_{2}=\mu\chi_{\Theta\setminus A}/(1-t)$, we can write $\mu$ as a convex combination
@@
The following lemma is contained in the proof of [18, Theorem 3.2]. It allows to reduce a minimization problem over compositional functions to a sequence of interpolation problems with respect to a generic minimizer. Since we found it of independent interest, we thought to emphasize it in a separate lemma.
Let $L$ be a positive integer. Take a set $\mathcal{X}_{0}$ and Banach spaces $\mathcal{X}_{1},\dots,\mathcal{X}_{L+1}$. For each $\ell=0,\ldots,L$, fix a Banach space $\mathcal{H}_{\ell}$ of functions from $\mathcal{X}_{\ell}$ to $\mathcal{X}_{\ell+1}$. To every $f{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=f_{0}\oplus\cdots\oplus f_{L}}\in\bigoplus_{\ell=0}^{L}\mathcal{H}_{\ell}$, recall that $f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}$ is defined as
$f^{\text{deep}}=f_{L}\circ\cdots\circ f_{0}\ .$
@@
[…]
@@
Assume that (17) has a solution $f^{*}=\oplus f_{\ell}^{*}$, and denote $x_{i}^{(0)}=x_{i}$ and $x_{i}^{(\ell+1)}=f^{*}_{\ell}(x_{i}^{(\ell)})$ for $\ell=0,\dots,L$. Then there exists a minimizer $\widetilde{f}=\oplus_{\ell=0}^{L}\widetilde{f}_{\ell}$ of (17) such that for all $\ell=0,\dots,L$ the function $\widetilde{f}_{\ell}$ is the solution of
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\inf_{f_{\ell}\in\mathcal{H}_{\ell}}}\|f_{\ell}\|_{\mathcal{H}_{\ell}}\quad\text{subject to}\quad f_{\ell}(x^{(\ell)}_{i})=x^{(\ell+1)}_{i}\qquad i=1\dots,N\ .$ (18)
@@
[…]
The next lemmas are needed to establish continuity in the setting of our representer theorem 4.1. A continuity result akin to A.7 is also needed for the proof of [18, Theorem 3.2]. We remark that, in order for the argument to go thorough, joint continuity is required, while the proof in [18, Theorem 3.2] only establishes separate continuity. The final result remains valid since joint continuity holds true nevertheless. This can be seen as a special case of our A.7, where the last steps can be simplified in view of the fact that, in finite-dimensional spaces, weak and strong continuity coincide.
@@
$\phi(x)\mu=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta)\qquad x\in\mathcal{X}\qquad\mu\in\mathcal{M}(\Theta,\mathcal{Y})\ .$
Then, for all $x\in\mathcal{X}$, $\phi(x)$ is continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ endowed with the $*{}^{*}$ topology to $\mathcal{Y}$ endowed with the weak topology.
@@
Thus $\langle\phi(x)\cdot,y\rangle_{\mathcal{Y}}$ defines an element of the predual $\mathcal{C}_{0}(\Theta,\mathcal{Y})$, and hence it is $*{}^{*}$ continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ to $\mathbb{R}$. But since this is true for all $y\in\mathcal{Y}$, it is $*{}^{*}$ continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ to $\mathcal{Y}$ endowed with the weak topology. ∎
@@
Let $\Theta$ be a Polish space. For all $n\in\mathbb{N}$, let $\lambda_{n}\in\mathcal{M}(\Theta)$ and $f_{n}:\Theta\to\mathbb{R}$ satisfying the following conditions:
@@
    the sequence $(\lambda_{n})$ converges to some $\lambda\in\mathcal{M}(\Theta)$ with respect to the narrow topology, i.e.
    $\lim_{n\to+\infty}\int_{\Theta}\varphi(\theta){\rm d}\lambda_{n}(\theta)=\int_{\Theta}\varphi(\theta){\rm d}\lambda(\theta)\qquad\forall\varphi\in\mathcal{C}_{b}(\Theta)\ ;$
@@
    the function $f$ is $\lambda$-integrable;
@@
$=$ $\left|\int_{K}(f_{n}(\theta)-f(\theta)){\rm d}\lambda_{n}(\theta)\right|+\left|\int_{\Theta\setminus K}(f_{n}(\theta)-f(\theta)){\rm d}\lambda_{n}(\theta)\right|+\left|\int_{\Theta}(f_{n}(\theta)-f(\theta)){\rm d}\lambda(\theta)\right|$
$\leq$ $\sup_{K}|f_{n}-f||\lambda_{n}(K)|+\sup_{\Theta\setminus K}|f_{n}-f||\lambda_{n}(\Theta\setminus K)|+\left|\int_{\Theta}(f_{n}(\theta)-f(\theta)){\rm d}\lambda(\theta)\right|\ .$
The last term goes to zero by dominated convergence since $f_{n}-f$ is uniformly bounded. The first term goes to zero because $f_{n}\to f$ uniformly on $K$ and, since $(\lambda_{n})$ is convergent, $|\lambda_{n}(K)|\leq|\lambda_{n}(\Theta)|\leq\sup_{n}\|\lambda_{n}\|_{\operatorname{TV}}<\infty$. For the middle term, fix an arbitrary $\varepsilon>0$. Again, $f_{n}-f$ is uniformly bounded. Moreover, since $(\lambda_{n})$ is convergent and $\Theta$ is Polish, by the Prokhorov theorem [6, Theorem 8.6.2] there is a compact set $K_{\varepsilon}\subset\Theta$ such that $|\lambda_{n}(\Theta\setminus K_{\epsilon})|<\varepsilon$ for all $n$. By taking $K=K_{\varepsilon}$ we get
@@
Let $\mathcal{X}_{0}$ be a set, $\mathcal{X}_{1},\mathcal{X}_{2}$ separable Hilbert spaces, and $\Theta_{0},\Theta_{1}$ locally compact, second countable topological spaces. For $\ell=0,1$, let $\rho_{\ell}:\mathcal{X}_{\ell}\times\Theta_{\ell}\to\mathbb{R}$ such that $\rho_{\ell}(x,\cdot)\in\mathcal{C}_{0}(\Theta_{\ell})$ for all $x\in\mathcal{X}_{\ell}$, and define $\phi_{\ell}(x):\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\to\mathcal{X}_{\ell+1}$ by
$\phi_{\ell}(x)\mu=\int_{\Theta_{\ell}}\rho_{\ell}(x,\theta){\rm d}\mu(\theta)\qquad x\in\mathcal{X}_{\ell}\qquad\mu\in\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\ .$
@@
is jointly $*{}^{*}$ continuous from $B_{\mathcal{M}(\Theta_{0},\mathcal{X}_{1})}(r_{0})\times B_{\mathcal{M}(\Theta_{1},\mathcal{X}_{2})}(r_{1})$ to $\mathcal{X}_{2}$ endowed with the weak topology.
@@
By the Banach–Alaoglu theorem, the product $B=B_{\mathcal{M}(\Theta_{0},\mathcal{X}_{1})}(r_{0})\times B_{\mathcal{M}(\Theta_{1},\mathcal{X}_{2})}(r_{1})$ is compact. Moreover, since $\mathcal{X}_{\ell+1}$ ($\ell=0,1$) is separable, so is $\mathcal{C}_{0}(\Theta_{\ell},\mathcal{X}_{\ell+1})$ by the Stone–Weierstrass theorem. Also, $\mathcal{X}_{\ell+1}$ is Hilbert, hence $\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})=\mathcal{C}_{0}(\Theta_{\ell},\mathcal{X}_{\ell+1})^{\prime}$ by the Riesz representation theorem. Therefore, $B$ is metrizable [20, Theorem 3.16]. Thus, it is enough to prove the ($*{}^{*}$-weak) sequential continuity of $\Gamma_{x}$.
To this end, let $(\mu_{n},\nu_{n})\to(\mu,\nu)$ ($*{}^{*}$). We want to show that $\Gamma_{x}(\mu_{n},\nu_{n})\to\Gamma_{x}(\mu,\nu)$ (weakly). We have
@@
The second term goes to zero by A.5. Let us call $\mathcal{I}$ the first term, and let $z_{n}=\phi_{0}(x)\mu_{n}$, $z=\phi_{0}(x)\mu$. For all $y\in\mathcal{X}_{2}$, by assumption (19) we have
@@
where $[\nu_{n}]_{y}(E)=\langle\nu_{n}(E),y\rangle_{\mathcal{X}_{2}}$ for all $E\in\mathcal{B}(\Theta_{1})$. We want to apply A.6 with $f_{n}(\theta)=|\langle z_{n}-z,g(\theta)\rangle_{\mathcal{X}_{1}}|$ and $\lambda_{n}=|\beta||[\nu_{n}]_{y}|$, so to conclude that $|\langle\mathcal{I},y\rangle_{\mathcal{X}_{2}}|\to 0$.
@@
where $\sup_{n}\|z_{n}-z\|_{\mathcal{X}_{1}}<\infty$ because $(z_{n}-z)$ is convergent, and $\|g\|_{\infty}<\infty$ by assumption. Next, we show that $(\lambda_{n})$ converges pointwise on $\mathcal{C}_{b}(\Theta_{1})$. Let $\lambda=|\beta||[\nu]_{y}|$ where $[\nu]_{y}(E)=\langle\nu(E),y\rangle_{\mathcal{X}_{2}}$ for all $E\in\mathcal{B}(\Theta_{1})$. Then, for every $h\in\mathcal{C}_{b}(\Theta_{1})$ we have $h|\beta|\in\mathcal{C}_{0}(\Theta_{1})$, and hence, since $[\nu_{n}]_{y}\to[\nu]_{y}$ pointwise on $\mathcal{C}_{0}(\Theta_{1})$,
@@
Finally, we show that $f_{n}\to 0$ uniformly on compact sets. Let $K\subset\Theta_{1}$ be compact, and fix an arbitrary $\varepsilon>0$. Since $g$ is continuous, $g(K)$ is compact in $\mathcal{X}_{1}$, and thus it can be covered by a finite number of closed balls. Let $w_{1},\dots,w_{q}\in\mathcal{X}_{1}$ be the centers of such balls, and define $P:\mathcal{X}_{1}\to\mathcal{X}_{1}$ as the projection onto $\operatorname{span}\{w_{1},\dots,w_{q}\}$. Then $\sup_{w\in g(K)}\|w-Pw\|_{\mathcal{X}_{!}}\leq\varepsilon$. Hence, for every $\theta\in K$ there is $w\in P\mathcal{X}_{1}$ such that $\|g(\theta)-w\|_{\mathcal{X}_{1}}\leq\varepsilon$. Thus, we have
@@
All the assumptions of A.6 are therefore satisfied, and its application concludes the proof. ∎

### v2

@@ 7. Conclusions and future work
Studying function spaces defined by neural networks provides a natural way to understand their properties. Recently, reproducing kernel Banach spaces have emerged has a useful concept to study shallow networks.
@@
In this paper, we take a step towards more complex architectures considering deep networks. We allow for a wide class of activation functions and remove unnecessary low rank constraints. Our main contributions are defining classes of neural RKBS obtained composing vector-valued RKBS and deriving corresponding representer theorems borrowing ideas from [22].
@@
Future developments include considering more structured architectures, for example convolutional networks, as well as investigating the statistical and computational properties of neural RKBS. Moreover, finer characterizations of the Banach structure could be obtained using the specific form of the activation function and the functional properties that this induces (see [22] for the ReLU).
Acknowledgments
L. R. acknowledges the financial support of the European Research Council (grant SLING 819789), the European Commission (Horizon Europe grant ELIAS 101120237), the US Air Force Office of Scientific Research (FA8655-22-1-7034), the Ministry of Education, University and Research (FARE grant ML4IP R205T7J2KP; grant BAC FAIR PE00000013 funded by the EU - NGEU) and the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF-1231216. This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains. The research by E.D.V. and L. R. has been supported by the MIUR grant PRIN 202244A7YL. The research by E.D.V. has been supported by the MIUR Excellence Department Project awarded to Dipartimento di Matematica, Università di Genova, CUP D33C23001110001. E.D.V. is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains. This work was partially supported by the MUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C18000100006. S.V. also acknowledges financial support from the MUR 2022 PRIN project GRAFIA, project code 202284Z9E4.
@@ Appendix A Sparse solutions to finite-dimensional variational problems
Let $Q$ be a convex subset of a locally convex space. A point $q\in Q$ is called extremal if $Q\setminus\{q\}$ is convex, that is, there do not exist $\in Q, prp\neq r$, such that q=tp+(1t)rq=tp+(1-t)r for some $t\in(0,1)$. We denote the set of extremal points of $Q$ by $\operatorname{Ext}(Q)$.
@@
˜A.2 is a simplified version of [7, Theorem 3.3], where $G$ is only assumed to be a seminorm.
In view of ˜A.2, in order to determine the form of sparse solutions we need to characterize the set of extremal point of the unit ball in $U$. In our construction of integral RKBS, $U$ is the space of vector measures with total variation norm. The following result provides the desired characterization for our case. It appears in [33] assuming that $\Theta$ is compact, but the proof works analogously when $\Theta$ is locally compact. We report the proof for the reader’s convenience.
Lemma A.3 ([33, Theorem 2]).
@@
Let us denote $E=\{y\cdot\delta_{\theta}:y\in\operatorname{Ext}(B_{\mathcal{Y}}(1)),\theta\in\Theta\}$. We start showing that $\operatorname{Ext}(B_{\mathcal{M}}(1))\subseteq E$. Suppose that $\mu\in\operatorname{Ext}(B_{\mathcal{M}}(1))$ but $\mu\neq y\cdot\delta_{\theta}$ for any $y\in\operatorname{Ext}(B_{\mathcal{Y}}(1))$ and $\theta\in\Theta$. Then $|\mu|\neq\delta_{\theta}$ for any $\theta\in\Theta$, and there is $A\in\mathcal{B}(\Theta)$ such that $0<|\mu|(A)<1$. Denote by $\chi_{A}$ the indicator function on $A$. Then, setting $t=|\mu|(A)$, $\mu_{1}=\mu\chi_{A}/t$ and $\mu_{2}=\mu\chi_{\Theta\setminus A}/(1-t)$, we can write $\mu$ as a convex combination
@@
The following lemma is contained in the proof of [22, Theorem 3.2]. It allows to reduce a minimization problem over compositional functions to a sequence of interpolation problems with respect to a generic minimizer. Since we found it of independent interest, we thought to emphasize it in a separate lemma.
Let $L$ be a positive integer. Take a set $\mathcal{X}_{0}$ and Banach spaces $\mathcal{X}_{1},\dots,\mathcal{X}_{L+1}$. For each $\ell=0,\ldots,L$, fix a Banach space $\mathcal{H}_{\ell}$ of functions from $\mathcal{X}_{\ell}$ to $\mathcal{X}_{\ell+1}$. To every $f{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=f_{0}\oplus\cdots\oplus f_{L}}\in\bigoplus_{\ell=0}^{L}\mathcal{H}_{\ell}$, recall that $f^{\rm deep}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}$ is defined as
$f^{\rm deep}=f_{L}\circ\cdots\circ f_{0}.$
@@
$\mathcal{A}:\bigoplus_{\ell=0}^{L}\mathcal{H}_{\ell}\to\mathbb{R}^{N},\qquad\mathcal{A}(f)_{i}=f^{\rm deep}(x_{i}).$
@@
Assume that (31) has a solution $f^{*}=\oplus_{\ell=0}^{L}f_{\ell}^{*}$, and denote $x_{i}^{(0)}=x_{i}$ and $x_{i}^{(\ell+1)}=f^{*}_{\ell}(x_{i}^{(\ell)})$ for $\ell=0,\dots,L$. Then there exists a minimizer $\widetilde{f}=\oplus_{\ell=0}^{L}\widetilde{f}_{\ell}$ of (31) such that for all $\ell=0,\dots,L$ the function $\widetilde{f}_{\ell}$ is the solution of
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\inf_{f_{\ell}\in\mathcal{H}_{\ell}}}\|f_{\ell}\|_{\mathcal{H}_{\ell}}\quad\text{subject to}\quad f_{\ell}(x^{(\ell)}_{i})=x^{(\ell+1)}_{i},\qquad i=1\dots,N,$ (32)
@@
But since $f^{*}$ is a minimizer, so is $\widetilde{f}$ and $\|\widetilde{f}_{\ell}\|_{\mathcal{H}_{\ell}}=\|f^{*}_{\ell}\|_{\mathcal{H}_{\ell}}$. ∎
Remark A.5.
We stress that the set $\{x_{i}^{(\ell)}:\ell=1,\ldots,L+1,i=1,\ldots,N\}$ defining the constraints in the problem (32) depends on the choice of the minimizer $f^{*}=\oplus_{\ell=0}^{L}f_{\ell}$ of the problem (31), so that there is some freedom to choose the points.
The next lemmas are needed to establish continuity in the setting of our representer theorem 5.3. A continuity result akin to ˜A.8 is also needed for the proof of [22, Theorem 3.2]. We remark that, in order for the argument to go thorough, joint continuity is required, while the proof in [22, Theorem 3.2] only establishes separate continuity. The final result remains valid since joint continuity holds true nevertheless. This can be seen as a special case of our ˜A.8, where the last steps can be simplified in view of the fact that, in finite-dimensional spaces, weak and strong continuity coincide.
@@
$\phi(x)\mu=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta),\qquad x\in\mathcal{X},\quad\mu\in\mathcal{M}(\Theta,\mathcal{Y}).$
Then, for all $x\in\mathcal{X}$, $\phi(x)$ is continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ endowed with the weak topology to $\mathcal{Y}$ endowed with the weak topology.
@@
Thus $\langle\phi(x)\cdot,y\rangle_{\mathcal{Y}}$ defines an element of the predual $\mathcal{C}_{0}(\Theta,\mathcal{Y})$, and hence it is weakly continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ to $\mathbb{R}$. But since this is true for all $y\in\mathcal{Y}$, it is weakly continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ to $\mathcal{Y}$ endowed with the weak topology. ∎
@@
Let $\Theta$ be a Polish space. For all $n\in\mathbb{N}$, let $\lambda_{n}\in\mathcal{M}(\Theta)$ and $f_{n}:\Theta\to\mathbb{R}$ continuous functions satisfying the following conditions:
@@
    the sequence $(\lambda_{n})_{n}$ converges to some $\lambda\in\mathcal{M}(\Theta)$ with respect to the narrow topology, i.e.
    $\lim_{n\to+\infty}\int_{\Theta}\varphi(\theta){\rm d}\lambda_{n}(\theta)=\int_{\Theta}\varphi(\theta){\rm d}\lambda(\theta),\qquad\forall\varphi\in\mathcal{C}_{b}(\Theta);$
@@
    the function $f$ is $\lambda$-integrable.
@@
$\leq$ $\left|\int_{K}(f_{n}(\theta)-f(\theta)){\rm d}\lambda_{n}(\theta)\right|+\left|\int_{\Theta\setminus K}(f_{n}(\theta)-f(\theta)){\rm d}\lambda_{n}(\theta)\right|+\left|\int_{\Theta}(f_{n}(\theta)-f(\theta)){\rm d}\lambda(\theta)\right|$
$+$ $\left|\int_{\Theta}f_{n}(\theta){\rm d}\lambda(\theta)-\int_{\Theta}f(\theta){\rm d}\lambda_{n}(\theta)\right|$
$\leq$ $\sup_{K}|f_{n}-f||\lambda_{n}|(K)+\sup_{\Theta\setminus K}|f_{n}-f||\lambda_{n}|(\Theta\setminus K)+\left|\int_{\Theta}(f_{n}(\theta)-f(\theta)){\rm d}\lambda(\theta)\right|$
$+$ $\left|\int_{\Theta}f_{n}(\theta){\rm d}\lambda(\theta)-\int_{\Theta}f(\theta){\rm d}\lambda_{n}(\theta)\right|.$
The first term goes to zero because $f_{n}\to f$ uniformly on $K$ and, since $(\lambda_{n})_{n}$ is convergent, $|\lambda_{n}|(K)\leq|\lambda_{n}|(\Theta)\leq\sup_{n}\|\lambda_{n}\|_{\operatorname{TV}}<\infty$. The third term goes to zero by dominated convergence since $(f_{n}-f)_{n}$ is uniformly bounded. The last term goes to zero since both terms converge to the integral $\int_{\Theta}f(\theta){\rm d}\lambda(\theta)$ as $n\to\infty$: the first term by the dominated convergence theorem since $(f_{n})_{n}$ is uniformly bounded, and the second term by hypothesis $(iv)$ since $f\in C_{b}(\Theta)$ as uniform limit on compact sets of a uniformly bounded sequence of continuous functions. For the second term, fix an arbitrary $\varepsilon>0$. Again, $(f_{n}-f)_{n}$ is uniformly bounded. Moreover, since $(\lambda_{n})_{n}$ is convergent and $\Theta$ is Polish, by the Prokhorov theorem [6, Theorem 8.6.2] there is a compact set $K_{\varepsilon}\subset\Theta$ such that $|\lambda_{n}|(\Theta\setminus K_{\epsilon})<\varepsilon$ for all $n$. By taking $K=K_{\varepsilon}$ we get
@@
Let $\mathcal{X}_{0}$ be a set, $\mathcal{X}_{1},\mathcal{X}_{2}$ separable Hilbert spaces, and $\Theta_{0},\Theta_{1}$ locally compact, second countable topological spaces. For $\ell=0,1$, let $\rho_{\ell}:\mathcal{X}_{\ell}\times\Theta_{\ell}\to\mathbb{R}$ be such that $\rho_{\ell}(x,\cdot)\in\mathcal{C}_{0}(\Theta_{\ell})$ for all $x\in\mathcal{X}_{\ell}$, and define $\phi_{\ell}(x):\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\to\mathcal{X}_{\ell+1}$ by
$\phi_{\ell}(x)\mu=\int_{\Theta_{\ell}}\rho_{\ell}(x,\theta){\rm d}\mu(\theta),\qquad x\in\mathcal{X}_{\ell},\qquad\mu\in\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1}).$
@@
is jointly weakly continuous from $B_{\mathcal{M}(\Theta_{0},\mathcal{X}_{1})}(r_{0})\times B_{\mathcal{M}(\Theta_{1},\mathcal{X}_{2})}(r_{1})$ to $\mathcal{X}_{2}$ endowed with the weak topology.
@@
By the Banach–Alaoglu theorem, the product $B=B_{\mathcal{M}(\Theta_{0},\mathcal{X}_{1})}(r_{0})\times B_{\mathcal{M}(\Theta_{1},\mathcal{X}_{2})}(r_{1})$ is compact. Moreover, since $\mathcal{X}_{\ell+1}$ ($\ell=0,1$) is separable, so is $\mathcal{C}_{0}(\Theta_{\ell},\mathcal{X}_{\ell+1})$ by the Stone–Weierstrass theorem. Also, $\mathcal{X}_{\ell+1}$ is Hilbert, hence $\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})=\mathcal{C}_{0}(\Theta_{\ell},\mathcal{X}_{\ell+1})^{\prime}$ by the Riesz representation theorem. Therefore, $B$ is metrizable [24, Theorem 3.16]. Thus, it is enough to prove the (weak-weak) sequential continuity of $\Gamma_{x}$.
To this end, let $(\mu_{n},\nu_{n})\to(\mu,\nu)$ (weakly). We want to show that $\Gamma_{x}(\mu_{n},\nu_{n})\to\Gamma_{x}(\mu,\nu)$ (weakly). We have
@@
The second term goes to zero by ˜A.6. Let us call $\mathcal{I}$ the first term, and let $z_{n}=\phi_{0}(x)\mu_{n}$, $z=\phi_{0}(x)\mu$. For all $y\in\mathcal{X}_{2}$, by assumption (33) we have
@@
where $\nu_{n}]_{y}(E)=\langle\nu_{n}(E),y\rangle_{\mathcal{X}_{2}}$ for all $E\in\mathcal{B}(\Theta_{1})$. We want to apply ˜A.7 with $f_{n}(\theta)=|\langle z_{n}-z,g(\theta)\rangle_{\mathcal{X}_{1}}|$ and $\lambda_{n}=|\beta||[\nu_{n}]_{y}|$, so to conclude that $|\langle\mathcal{I},y\rangle_{\mathcal{X}_{2}}|\to 0$.
@@
where $\sup_{n}\|z_{n}-z\|_{\mathcal{X}_{1}}<\infty$ because $(z_{n}-z)$ is convergent, and $\|g\|_{\infty}<\infty$ by assumption. Next, we show that $(\lambda_{n})$ converges pointwise on $\mathcal{C}_{b}(\Theta_{1})$. Let $\lambda=|\beta||[\nu]_{y}|$ where $\nu]_{y}(E)=\langle\nu(E),y\rangle_{\mathcal{X}_{2}}$ for all $E\in\mathcal{B}(\Theta_{1})$. Then, for every $h\in\mathcal{C}_{b}(\Theta_{1})$ we have $h|\beta|\in\mathcal{C}_{0}(\Theta_{1})$, and hence, since $\nu_{n}]_{y}\to[\nu]_{y}$ pointwise on $\mathcal{C}_{0}(\Theta_{1})$,
@@
Finally, we show that $f_{n}\to 0$ uniformly on compact sets. Let $K\subset\Theta_{1}$ be compact, and fix an arbitrary $\varepsilon>0$. Since $g$ is continuous, $g(K)$ is compact in $\mathcal{X}_{1}$, and thus it can be covered by a finite number of closed balls of radius $\varepsilon$. Let $w_{1},\dots,w_{q}\in\mathcal{X}_{1}$ be the centers of such balls, and define $P:\mathcal{X}_{1}\to\mathcal{X}_{1}$ as the projection onto $\operatorname{span}\{w_{1},\dots,w_{q}\}$. Then $\sup_{w\in g(K)}\|w-Pw\|_{\mathcal{X}_{1}}\leq\varepsilon$. Hence, for every $\theta\in K$ there is $w\in P\mathcal{X}_{1}$ such that $\|g(\theta)-w\|_{\mathcal{X}_{1}}\leq\varepsilon$. Thus, we have
@@
All the assumptions of ˜A.7 are therefore satisfied, and its application concludes the proof. ∎
Remark A.9.
In the framework of deep integral RKBS, the above lemma states that the evaluation functional at a fixed $x\in\mathcal{X}_{0}$
$(\mu,\nu)\mapsto f_{\nu}(f_{\mu}(x))$
is jointly weakly continuous from $B_{\mathcal{M}(\Theta_{0},\mathcal{X}_{1})}(r_{0})\times B_{\mathcal{M}(\Theta_{1},\mathcal{X}_{2})}(r_{1})$ to $\mathcal{X}_{2}$ endowed with the weak topology. The proof for $L=2$ can be easily extended to cover the case $L>2$.